OpenStudy (darkbluechocobo):

Help with Finding the angle between 2 lines

OpenStudy (darkbluechocobo):

Find the angle between the following two lines: y = x x − 2y = 0

OpenStudy (darkbluechocobo):

y=1/2x

OpenStudy (darkbluechocobo):

1/2-1=-1/2 1/2*1+1=3/2

OpenStudy (darkbluechocobo):

\[\tan^{-1}(\frac{ -1/2 }{ 3/2 })\]

OpenStudy (darkbluechocobo):

Theta=18.4 @freckles does this look right?

OpenStudy (darkbluechocobo):

because its \[\left| \frac{ -1/2 }{ 3/2 } \right|\]

OpenStudy (darkbluechocobo):

@Loser66 does this look right?

OpenStudy (freckles):

we will do it your way since it is easier :p ok let's look the slopes \[m_1=1, m_2=\frac{1}{2} \\ m_1-m_2=\frac{1}{2} \\ 1+m_1 \cdot m_2 =1+1 \cdot \frac{1}{2}=\frac{3}{2} \\ \text{ so yes you have } |\frac{\frac{1}{2}}{\frac{3}{2}}|=|\frac{1}{2} \cdot \frac{2}{3}|\]

OpenStudy (freckles):

then you do arctan( of that )

OpenStudy (darkbluechocobo):

can confirm 18.4 is what i got so I did it right this time

OpenStudy (freckles):

yea to the nearest 10th yeah

OpenStudy (darkbluechocobo):

Thank you :D

OpenStudy (freckles):

np